# General Phrase Debiaser: Debiasing Masked Language Models at a Multi-Token Level

### v1

@@ 1 Introduction
Recently, masked language models (MLMs) [1, 2, 3, 4, 5] are employed in both traditional tasks like text classification and diverse multimodal tasks when combined with models like image generators [6, 7]. We aim to develop MLMs with minimal human biases, even when the pretraining data unavoidably contains these biases. However, correcting implicit biases in pretrained MLMs can be very challenging, especially considering the high cost of retraining models from scratch.
@@ 3.1 Evaluation Data And Details
Implementation Details: In step 2 of Fig.1, we use $S_{unweighted}$ which has 624 phrases. The maximum biased prompt length $PL$ is 5 and beam search width $K$ is 100. We use the 5,000 highest frequency words in Wikipedia as the search space $\mathcal{V}$, to avoid noise in the vocabulary $\mathcal{V}$ and speed up the prompt search process. In step 3, we use $S_{weighted}$ with more than 500 phrases instead of $S_{unweighted}$ because $S_{weighted}$ takes into account the varying weights of different stereotypical phrases, resulting in better debiasing effects (as shown in Table 1). And we choose $\mathcal{C}^{*}$ (an extension of $\mathcal{C}$, derived from the gender word list in [9]) as attribute phrases to construct more fine-tuning data. All models are trained with AdamW [19] optimizer and early stopping strategy. Our experiments run on a single NVIDIA 3090Ti.

### v2

@@ 1 Introduction
Recently, masked language models (MLMs) [1, 2, 3, 4, 5, 6] are employed in both traditional tasks like text classification [7, 8, 9] and diverse multimodal tasks [10, 11] when combined with models like image generators [12, 13]. We aim to develop MLMs with minimal human biases, even when the pretraining data unavoidably contains these biases. However, correcting implicit biases in pretrained MLMs can be very challenging, especially considering the high cost of retraining models from scratch.
@@ 3.1 Evaluation Data And Details
Implementation Details: Hyperparameters play a critical role in final performance [25, 26, 27]. For completeness sake, we then introduce the hyperparameters we used in our study. In step 2 of Fig.1, we use $S_{unweighted}$ which has 624 phrases. The maximum biased prompt length $PL$ is 5 and beam search width $K$ is 100. We use the 5,000 highest frequency words in Wikipedia as the search space $\mathcal{V}$, to avoid noise in the vocabulary $\mathcal{V}$ and speed up the prompt search process. In step 3, we use $S_{weighted}$ with more than 500 phrases instead of $S_{unweighted}$ because $S_{weighted}$ takes into account the varying weights of different stereotypical phrases, resulting in better debiasing effects (as shown in Table 1). And we choose $\mathcal{C}^{*}$ (an extension of $\mathcal{C}$, derived from the gender word list in [15]) as attribute phrases to construct more fine-tuning data. All models are trained with AdamW [28] optimizer and early stopping strategy. Our experiments run on a single NVIDIA 3090Ti.